Use the digits to make two 5 digit numbers and subtract one from the other to get an answer of or close to 12345

Answer by Edwin McCravy(20064)

Here are all possible ways, computer generated:

27948 - 15603 = 12345
28704 - 16359 = 12345
28740 - 16395 = 12345
29685 - 17340 = 12345
38019 - 25674 = 12345
38064 - 25719 = 12345
38091 - 25746 = 12345
38109 - 25764 = 12345
39801 - 27456 = 12345
39810 - 27465 = 12345
39846 - 27501 = 12345
49125 - 36780 = 12345
49152 - 36807 = 12345
49170 - 36825 = 12345
49215 - 36870 = 12345
50169 - 37824 = 12345
50214 - 37869 = 12345
50241 - 37896 = 12345
50286 - 37941 = 12345
50619 - 38274 = 12345
51024 - 38679 = 12345
51042 - 38697 = 12345
51069 - 38724 = 12345
51807 - 39462 = 12345
52086 - 39741 = 12345
60258 - 47913 = 12345
60537 - 48192 = 12345
61275 - 48930 = 12345
61302 - 48957 = 12345
61320 - 48975 = 12345
61725 - 49380 = 12345
62058 - 49713 = 12345
62103 - 49758 = 12345
62130 - 49785 = 12345
62175 - 49830 = 12345
63129 - 50784 = 12345
63174 - 50829 = 12345
63192 - 50847 = 12345
63219 - 50874 = 12345
72498 - 60153 = 12345
72534 - 60189 = 12345
72543 - 60198 = 12345
74235 - 61890 = 12345
74253 - 61908 = 12345
74280 - 61935 = 12345
74325 - 61980 = 12345
82659 - 70314 = 12345
83604 - 71259 = 12345
83640 - 71295 = 12345
84396 - 72051 = 12345
93765 - 81420 = 12345
94701 - 82356 = 12345
94710 - 82365 = 12345
95412 - 83067 = 12345
95421 - 83076 = 12345
96357 - 84012 = 12345

Edwin